If you have 44 dimes and quarters with a total of $8.60, how many dimes are there?
Found 2 solutions by Alan3354, Edwin McCravy:
Answer by Alan3354(69443)

If you have 44 dimes and quarters with a total of $8.60, how many dimes are there?
--------
d + q = 44
10d + 25q = 860
================
Or,
If it's 44 dimes, it's worth $4.40
Each dime replaced by a quarter add $0.15
---------
860 - 440 = 420
420/15 = 28
--> 28 quarters

Answer by Edwin McCravy(20059)

If you have 44 dimes and quarters with a total of $8.60, how many dimes are there?
D = number of dimes.
44-D = number of quarters

$.10D + $.25(44-D) = $8.60

 .10D +  .25(44-D) =  8.60
  10D +   25(44-D) =   860
  
You finish solving that and get D = 16.

D = 16 = number of dimes 
44-D = 44-16 = 28 = number of quarters

Checking: 16 dimes    is $1.60
          28 quarters is $7.00
------------------------------
                total is $8.60

Edwin